Help please somebody
I'm having a problem with the following question
The distance of a body from a point is given by s=xsin(wt+alpha). Show that the velocity and acceleration are given by v=x cos(wt+alpha)and a=-w^2s
Point me in the right direction please
Cheers Paul

Answer by khwang(438)

Given distance s=xsin(wt+alpha).
Show that the velocity and acceleration are
given by v=x cos(wt+alpha)and a=-w^2s

v = ds/dt (derivative of s wrt t)
= x w cos (wt+alpha)(by chain rule)
a = dv/st = -x w^2 sin(wt+alpha)= -w^2 s
something wrong with your typing about v.

Kenny